a black hole has a mass that is 10,000 times the mass of the sun. what is its schwarzschild radius? show all work for full credit.
Answer #1

so here mass = 104 * solar mass .... r = 2.95 * 104  km
Answer #2

Mass of sun =1.98*10^30kg

The radius is given by

rs =2GM/c^2 =1.48*10^-27*1000*1.98*10^30=2930*10^6km
Answer #3

Given black hole mass M = 10000 times the mass of the sun

mass of sun = 2*1030 kg

so here M = 2 * 1030 * 10000

Schwardzschild radius rs = 2 * G * M / C2

where G = gravitational constant

C = speed of the light

M = mass of the black hole

the value of 2G/ C2 = 1.48 * 10-27 m/kg

rs =( 2 * 6.67*10^-11 * 2* 10000) / ( 3*108)2 = 1.48 * 10-27 * 2 * 1030 * 104

rs = 2.96 * 107 m